OpenStudy (anonymous):

A compound consists of 3.47 grams of oxygen and 1.52 grams of nitrogen. What is the empirical formula of this compound? a. NO b. N2O2 c. N2O d. NO2

OpenStudy (mertsj):

Did you change the grams to moles so you know the molar relationship?

OpenStudy (mertsj):

3.47 g O = .2166 mole O atoms 1.52 g N = .1096 mole N atoms Ratio : \[\frac{.2166 mole O atoms}{.1096 mole N atoms}=\frac{1.995}{1}=\frac{2 O}{1 N}\]

OpenStudy (mertsj):

Empirical Formula: \[NO _{2}\]